when 1000cm3 of a 0.1M of NaOH solution is neutralized by 0.1M H2SO4 what mass of NaSO4 would be formed


1
Expert's answer
2021-09-16T02:01:01-0400

2NaOH+h2so4=na2so4+2H2O

1000cm3=1000ml

v(naoh)=1000/22400=0.0045mol

v(na2so4)=0.0022mol

m(na2so4)=142*0.0022=3.17g